OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit

OpenStudy (anonymous):

\[a_n= \frac{ (-1)^n * n }{ n^2+6 }\]

OpenStudy (anonymous):

I figured the (-1)^n = divergent.... but what happens when it is over infinity?

OpenStudy (anonymous):

Well it looks like the numerator alternates back and forth between -n and n. Do you agree?

OpenStudy (anonymous):

yes, that is why i said divergent for that one

OpenStudy (anonymous):

Good thinking. However, the denominator is of a higher degree, so which will grow quicker, the numerator or the denominator?

OpenStudy (anonymous):

\[\lim n--> \inf a_n= \frac{ (-1)^n }{ 1+6/n }\]

OpenStudy (anonymous):

im thinking the numerator will

OpenStudy (anonymous):

oh wait, thats just one...so denominator will

OpenStudy (anonymous):

Well, if the numerator grows with n and the denominator grows with n^2, which one gets bigger faster?

OpenStudy (anonymous):

the denominator. but, doesnt the divergent numerator play a role? Or do we ignore that

OpenStudy (anonymous):

Well, I'm thinking that, since the denominator grows more quickly than the numerator, the function tends toward zero. The numerator means that the function varies from positive value to negative value to positive value to negative value, etc., all the while getting closer to zero. Make sense?

OpenStudy (anonymous):

yes, this is what confuses me so

OpenStudy (anonymous):

so, im thinking convergent at zero

OpenStudy (anonymous):

or just divergent ... cuz the numerator still throws me off

OpenStudy (anonymous):

Me too. Perhaps like this

OpenStudy (anonymous):

i see what youre saying, either way, it still gets lower over time

OpenStudy (anonymous):

Right on.

OpenStudy (anonymous):

sweet, thanks

OpenStudy (anonymous):

You're welcome